# library.bib

### v1

@@ Discussion
In agreement with our prediction of the optimal spatial arrangement for maximum information transfer, spider monkeys seem to be balancing the intersection/union ratio of their core areas, or the extent of redundant and unique knowledge that an interaction would provide to a given set. Optimality approaches to collective behaviour are scarce, due to the difficulty in identifying the currency and level of optimization [marshall2024aims]. Here we have informed our representation of higher-order interactions as simplicial complexes with this theoretical optimal prediction of a convex decreasing function of the size of the subset. This optimality argument, developed in detail in the Supplementary Information, assumes that the intersections between individuals comprising each simplex provide an opportunity to share unique information that each one (or subsets of them) possess. In other words, our prediction is simply based upon the balance between redundancy and uniqueness in the use of space by any given set of individuals. This balance is not what one would propose for a central place forager (e.g. hooded crows, Corvus corone cornix [sonerud2001ignorant]) which would only share a small portion of redundant areas for sets of any size, as the opportunities for information sharing would only take place in a small, regularly used central area within their range (e.g. their roost). Our optimality argument also rests on the assumption that each group member benefits from information pooling in a way that surpasses the costs of feeding competition [clark1984foraging] and the alternative benefit of gathering information individually [valone2002public].
@@
While other, alternative models could possibly explain the observed patterns, our model simply assumes a balance between the opportunities to coincide in overlapping areas and the maximum amount of knowledge that can be shared about areas outside of those overlaps, applicable to any number of individual areas. For example, the most redundant interactions could be due to individuals cohesively exploring common areas for other, social reasons. Conversely, values below the predicted line are rare: sets of individual ranges do not seem less redundant than would be predicted based on our optimality argument. This is consistent with previous findings on the same species [ramos2020collective] indicating that repulsion between individuals is less relevant than attraction as a mechanism behind fission-fusion dynamics.
@@
Particularly for smaller subsets, between 2 and 6 individuals, the value of the intersection/union ratio falls steeply with $n$, both for the optimal $w^{*}$ and the observed values of $w$. This suggests that the balance between redundant and unique knowledge is more sensitive to the addition of new areas, contrary to the situation with subsets of 10 individuals or larger. It is interesting that the size of those subsets falls within the range for which we observed the largest variation in simplicial centrality. This metric increases sigmoidally as simplices increase in size from 1 to 5 or 6. Simplices representing subgroups of that range of sizes could thus be key in maintaining updated levels of foraging information, as they are not only central in the simplicial complex (and therefore tend to participate in more simplices) but have their intersection/union ratio poised at the most sensitive range with respect to the addition of new individuals (and are therefore the most dynamic ones in terms of the redundant and unique knowledge they possess).
@@
Sueur et al. [sueur2011collective] predicted that broker individuals, who connect different clusters of the association network, are particularly important in fluid groups with high degrees of fission-fusion dynamics. Our simplicial degree centrality results show that subsets of increasingly many individuals participate in increasingly many spatial interactions with other simplices. The fact that we did not find a clear relationship between simplicial centrality and the age and sex of the adults, or their recent immigration status, suggests that rather than being an attribute of individuals (as in killer whales, for example: [brent2015ecological]), the asymmetry in foraging information is predominantly the result of spider monkeys’ high mobility and fluid grouping patterns. This had already been suggested by previous work on the factors that determine an individual’s decision towards fission or fusion [ramos2014unraveling] and is compatible with the idea that spider monkeys are collectively pooling their foraging information in a dynamic fashion [palacios2019uncovering]. Of course, many other factors related to individual age and sex classes shape the structure of dyadic association networks [ramos2009association], but here we are uncovering those aspects of fission-fusion dynamics that relate to foraging knowledge as a dynamic, distributed process which varies depending on all adult members’ knowledge about a dynamic environment. Future work analysing associations in subgroups as higher-order interactions will be useful to evaluate the relevance of foraging knowledge as we have uncovered it here as an additional influence on the structure of association networks.
@@
We did not find stronger evidence for complementarity in the dry compared to the wet seasons. While it appeared as though the simplicial complexes had more holes of various dimensions during the dry seasons, we were unable to detect statistically significant differences using a composite index of filtration complementarity. During the dry seasons, spider monkeys tend to rely on less abundant tree species and thus exploit a more heterogenous environment. During the wet season, the hyperabundance of Brosimum alicastrum would make it less valuable to share information about scarce sources of food. However, we did not find significant differences in the temporal variation of the biweekly fruit abundance between the seasons. While our sampled number of seasons is small, it is possible that rather than temporal variation, it is the spatial variation that is higher in the dry seasons than in the wet season. During the dry seasons, fruit is more scarce and dispersed, subgroups tend to be smaller and home ranges larger [smith2016seasonal]. It would be interesting to test the information content of different regions of the home range, using more accurate tree distribution and abundance data than was available for this study.
@@
More generally, finding resources in a heterogenous, dynamic environment can be seen as a problem of pattern recognition, which can be solved effectively by systems with some degree of distributed information processing. This capacity can be found in various natural systems, recently described as ‘liquid brains’ [sole2019liquid], ranging from social insect colonies to the immune system. These systems process information via interactions between mobile agents that interact only temporarily when they happen to be within reach. As an example, Vining et al. [vining2019mobility] modelled a set of simple agents that could only communicate with others within a short radius but were able to move about, thus increasing the number of agents with whom they could exchange information over time. This system was able to solve a global problem of density classification (converging to the binary state that was more common in the beginning of the run amongst all agents). Crucially, agents were able to solve the most difficult versions of the problem (where the difference between the majority and minority binary state was very small) only when moving at intermediate speeds. This allowed them to balance the frequency of interactions with others in their current neighbourhood and those in other locations. In other words, an optimal degree of mixing between agents led the group to solve more difficult problems effectively [vining2019mobility]. This optimal degree of mixing is analogous to the optimization of $w$ we used here: an optimal area of intersection between individual use areas allows monkeys to coincide to share information while exploring enough area on their own which can be shared. Unique information, gathered by individual members of the group or subsets of them, becomes complementary when monkeys coincide in redundant areas and can share what they uniquely know. The redundancy/uniqueness balance is also more generally related to that between exploitation, where more redundant information is gathered about the environment, and exploration, where more unique information is gathered [hills2015exploration, luppi2024information].
@@
Collective information processing can result from individual members sharing complementary information, as we have found here, or it may also include synergistic information [williams2010nonnegative], whereby two or more individual estimations of a target variable not only are complementary in their estimation but combine their information in such a way as to produce new knowledge [bettencourt2009rules]. An example would be if one subset of individuals would contribute the location of a food source and another subset the timing of the fruiting of that source. The latter could be done by individuals who happened to visit the same trees repeatedly, gathering unique information about their fruiting status (as has been observed in other high fission-fusion species like the chimpanzee (Pan troglodytes; [janmaat2016spatio]). The resulting, combined knowledge by both subsets of individuals would be synergistic in the sense of allowing all of them to exploit the food source according to its location and timing. Further empirical and modelling work could explore these ideas and search for synergy in collective information processing about the foraging environment in this and other species.
@@
Using a theoretical prediction of the balance between the use of redundant and unique areas, as well as simplicial complexes to represent higher-order interactions, we have been able to show complementarity in information sharing, a hallmark of collective information processing [bettencourt2009rules, garg2025causal]. This work exemplifies the usefulness of higher-order methods, as we have been able to uncover details in the spatial structure which could not be found in a dyadic representation of the same structure. While these methods come with greater complexity over the dyadic approach, they are extremely well suited to tackling problems surrounding collective behaviour. Spider monkeys with different knowledge of their foraging area contribute with different pieces of information to group-level knowledge, in such a way that the group as a whole can access more foraging trees in a dynamic landscape than any individual could on its own. This complementary information sharing in higher-order spatial networks would be a compelling example of collective intelligence in natural conditions [hutchins1995cognition].
@@ Methods
Ranging and subgroup composition data was collected by experienced observers between January 2012 and December 2017 in a habituated group of Geoffroy’s spider monkeys (Ateles geoffroyi) living around the Punta Laguna lake in the Otoch Ma’ax yetel Kooh protected area, in the Yucatan peninsula, Mexico. This group was the subject of a long-term, continuous study from 1997 to 2020 [ramos2018ecologia]. We defined dry seasons from November to April and wet seasons from May to October [smith2016seasonal]. The spider monkeys’ foraging environment differs crucially between the two seasons, due mainly to a very abundant species (Brosimum alicastrum) fruiting during the wet season [smith2016seasonal]. Therefore, fruit during the wet season is more abundant and homogeneously distributed, while during the dry season spider monkeys use more species that are less abundant overall and thus face a more heterogenous and uncertain environment [smith2016seasonal].
We tracked changes in fruit abundance by recording, every two weeks, the number of trees with fruit for each species in a phenological trail. This trail contained a sample of 10 individual trees from 16 species, including 12 which comprise > 50% of the monkey’s diet [ramos2001patterns, pinacho2010patrones]. We estimated the density of each of these species using data collected in 8 0.25ha plots (total of 2ha) and 48 100m x 2m line transects (total 0.96 ha) within the home range of the spider monkeys (see [bonilla2008forest] for more details about the vegetation sampling).
@@
We used individual locations for the estimation of individual core ranges, defined as the 60% utilization distribution generated with the adaptive Local Convex Hull (a-LoCoH) method developed by [getz2007locoh], using the T-LoCoH package [lyons2013home] for the R programming environment [getz2007locoh] with a constant value of 15000 for the a parameter for hull construction across seasons. The use of the 60% utilization distribution as a meaningful area is validated by previous work where we showed that it lied at the inflection point in the curve of isopleth values versus proportion of home range covered [ramos2013site]. Core ranges defined in this manner represent the area where individuals concentrated their regular movements, excluding more sporadically used areas. For more details on the calculation of individual core ranges, see [smith2016seasonal].
@@
We calculated the intersections and unions between individual core ranges for each season using the Spatstat package [baddeley2015spatial] for the R programming environment [rcoreteam2021R]. For a set $C=\{C_{1},\dots,C_{n}\}$ of $n$ individual core ranges, we used the ratio denoted as $w$ of the sizes of their intersection and their union:
@@
We used the fruit abundance data from the phenological trail, together with the diameter at breast height of the trees monitored and the density for each species in the study site, to estimate fruit abundance and its variability in each season. We used the following formulation, based on [smith2016seasonal]:
@@
where $\operatorname{IFA}f$ is the index of fruit abundance for the monitoring period $f$ based on the 16 species monitored, $\operatorname{Tf}_{i}$ is the proportion of trees with fruit of species $i$, $\operatorname{DBH}_{i}$ is the sum of the diameter at breast height over all the sampled trees with fruit of species $i$ during period $f$ and $\operatorname{Den}_{i}$ is the tree density for species $i$ in the study site. We considered the $\operatorname{IFA}_{f}$ values for all the monitoring periods in a season to obtain the mean and the coefficient of variation of the $\operatorname{IFA}_{f}$ per season.
@@
For each dataset (i.e., each season), we represented the entire space-sharing network of the study group as a simplicial complex $K$. The nodes of the network were all the adult individuals in the group during the particular season of the dataset (Table SI1). Since all core ranges coincided in some small common area (e.g. for sleep), we defined a maximal simplex size, $n_{\text{max}}$. This maximal size was needed because including interactions among more than $n_{\text{max}}$ individuals yielded a trivial simplicial complex structure (containing every possible simplex), and these larger interactions were less likely to be related to foraging activity. Then there exists a simplex connecting a subset of individuals if their corresponding relative core range intersection, $w$, is non-zero, and if the size of the subgroup is no more than $n_{\text{max}}$. Upon observation, the value of $n_{\text{max}}=6$ was the largest which consistently resulted in non-trivial structures across all seasons. We computed the Betti numbers of the simplicial complexes for all seasons.
@@
We investigated how the final simplicial complex, $K$, was formed when adding simplices in decreasing order of redundancy. Between subsets of the same size, we assumed a higher relative spatial overlap, quantified by $w$ corresponded to a more relevant interaction. What constituted a redundant spatial overlap for information transfer depends upon the number of individuals in the subset, and therefore we required a scaling of the $w$ values with their corresponding subset size in order to make fair comparisons. To this end, we computed the $w$ value from the spatial overlap of $n$ individuals which optimised the information transfer when these individuals had equal domain knowledge and utilised their core ranges uniformly, balancing the size of the unique areas they could share information about with the opportunity to share them when coinciding with others. As detailed in the Supplementary Information, for $n$ individuals this optimal overlap was computed as:
@@
This expression describes the amount of overlap required for a group of $n$ individuals to maximise their transfer of information under the given assumptions of uniformity and independence of movements. We then defined a parameter $\alpha$ to describe the redundancy of a given overlap. For the $i$-th non-zero spatial overlap in the data, consisting of $n(i)$ many individuals and with relative overlap value $w_{i}$, the corresponding $\alpha$ value was given as the solution to:
@@
where $f(\alpha)$ is some arbitrary positive and decreasing function. We picked $f(\alpha)=5-\alpha$ as a linear function such that the corresponding $\alpha$ value for each set was non-zero. We varied $\alpha$ over the interval $[0,5)$. Then, we defined a sequence of simplicial complexes $K_{0}\subseteq K_{1}\subseteq\dots\subseteq K_{N}=K$. Here, $K_{0}$ is the set of individuals in the study group (without any simplices connecting them) and each $K_{j}$ is a simplicial complex which adds the simplex corresponding to the $j$-th lowest $\alpha$ value to the complex $K_{j-1}$ (possibly adding additional lower-order simplices to satisfy downward closure). This sequence is called a filtered simplicial complex, and $\alpha$ is called the filtration parameter [Otter2017]. Simplicial complexes which are lower in the sequence will consist of more redundant interactions in terms of their information sharing.
We computed the Betti numbers for each simplicial complex in the sequence (the persistent homology) for all available years and seasons in order to examine how the higher-order structure varied with the redundancy of interactions. Analysis was performed using the Gudhi package for Topological Data Analysis, implemented in Python [gudhi2015urm].
@@
We quantified how much evidence of complementarity and information sharing there was in a filtered simplicial complex through the filtration complementarity index. Structures with greater evidence of complementarity and sharing would be those with a larger number of higher order features (i.e. larger Betti numbers of high order), which persisted for longer throughout the filtration. Simplicial complexes with holes of higher dimensions may be described as more complex and present more evidence of complementary information sharing among a wider variety of simplices. We therefore defined the filtration complementarity index for season $i$ as:
@@
Here, $B_{i}^{(d)}(\alpha)$ is the $d$-th Betti number at the filtration value of $\alpha$ in the $i$-th season, $n_{i}$ is the number of individuals in the data in this season, and $\langle\cdot\rangle_{[0,5)}$ is the average operator over the filtration interval $[0,5]$. We scaled the number of connected components by the number of individuals as to not give seasons with larger sample sizes a higher complexity score automatically (as the number of connected components is highly dependent upon $n_{i}$). We did not scale the higher-order features as we did not expect them to be (directly) impacted by sample size differences observed in our data. Higher-order features were given greater importance through the scaling by $d+1$. We also tested other definitions of the filtration complementary index, including non-weighted or which did not consider the connected components. All of them yielded similar results.
@@
We additionally studied which subgroups were the most central in the simplicial complex across years and seasons. For this, we implemented the maximal simplicial degree centrality measure as defined in [Serrano2020]. This is a generalisation of the standard degree centrality used in dyadic networks to simplicial complexes, designed so that the downward closure assumption does not produce biases towards higher or lower order simplices. Hence the simplices are not automatically classed as more or less central just for being part of a nested structure. We computed the centrality value for each simplex in the simplicial complex for each year and season at a filtration value of $\alpha=4$ (meaning we do not consider subgroups with more unique interactions). Other values of $\alpha$ yielded similar results.
@@
The individual core ranges analysed, as well as the data used for constructing the simplicial complexes are publicly available as part of the repository cited above.
[…]

### v3

@@ Discussion
In agreement with our prediction of the optimal spatial arrangement for maximum information transfer, spider monkeys seem to be balancing the intersection/union ratio of their core areas, or the extent of redundant and unique knowledge that an interaction would provide to a given set. Optimality approaches to collective behaviour are scarce, due to the difficulty in identifying the currency and level of optimization [34]. Here we have informed our representation of higher-order interactions as simplicial complexes with this theoretical optimal prediction of a convex decreasing function of the size of the subset. This optimality argument, developed in detail in the Supplementary Information, assumes that the intersections between individuals comprising each simplicial complex provide an opportunity to share unique information that each one (or subsets of them) possess. In other words, our prediction is simply based upon the balance between redundancy and uniqueness in the use of space by any given set of individuals. This balance is not what one would propose for a central place forager (e.g. hooded crows, Corvus corone cornix [8]) which would only share a small portion of redundant areas for sets of any size, as the opportunities for information sharing would only take place in a small, regularly used central area within their range (e.g. their roost). Our optimality argument also rests on the assumption that each group member benefits from information pooling in a way that surpasses the costs of feeding competition [4] and the alternative benefit of gathering information individually [6].
@@
While other, alternative models could possibly explain the observed patterns, ours simply assumes a balance between the opportunities to coincide in overlapping areas and the maximum amount of knowledge that can be shared about areas outside of those overlaps, applicable to any number of individual areas. For example, the most redundant interactions could be due to individuals cohesively exploring common areas for other, social reasons. Conversely, values below the predicted line are rare: sets of individual ranges do not seem less redundant than would be predicted based on our optimality argument. This is consistent with previous findings on the same species [35] indicating that repulsion between individuals is less relevant than attraction as a mechanism behind fission-fusion dynamics.
@@
Particularly for smaller subsets, between 2 and 6 individuals, the value of the intersection/union ratio falls steeply with $n$, both for the optimal $w^{*}$ and the observed values of $w$. This suggests that the balance between redundant and unique knowledge at this range of subset sizes is more sensitive to the addition of new areas, contrary to the situation with subsets of 10 individuals or larger. It is interesting that we found a similar pattern in the simplicial centrality analysis: this metric increases sigmoidally as simplices increase in size, with a maximum rate of increase in simplices of intermediate size (4 to 6 individuals) and then levelling off at larger simplex sizes of 7-9, depending on the season and the maximum simplex size considered (Figures SI5-9). From the overall patterns of simplicial centrality, we can speculate that simplices representing subsets of intermediate size could be key in maintaining updated levels of foraging information, as their centrality is more sensitive to simplex size and also have their intersection/union ratio poised at the most sensitive range with respect to the addition of new individuals (and are therefore the most dynamic ones in terms of the redundant and unique knowledge they possess).
@@
Sueur et al. [16] predicted that broker individuals, who connect different clusters of the association network, are particularly important in fluid groups with high degrees of fission-fusion dynamics. Our simplicial centrality results show that subsets of increasingly many individuals participate in increasingly many spatial interactions with other simplices. The fact that we did not find a clear relationship between simplicial centrality and the age and sex of the adults, or their recent immigration status, suggests that rather than being an attribute of individuals (as in killer whales, for example: [36]), the asymmetry in foraging information is predominantly the result of spider monkeys’ high mobility and fluid grouping patterns. This had already been suggested by previous work on the factors that determine an individual’s decision towards fission or fusion [19] and is compatible with the idea that spider monkeys are collectively pooling their foraging information in a dynamic fashion [11]. Of course, many other factors related to individual age and sex classes shape the structure of dyadic association networks [37], but here we are uncovering those aspects of fission-fusion dynamics that relate to foraging knowledge as a dynamic, distributed process which varies depending on all adult members’ knowledge about a dynamic environment. Future work analysing associations in subgroups as higher-order interactions will be useful to evaluate the relevance of foraging knowledge as we have uncovered it here as an additional influence on the structure of association networks.
@@
We did not find stronger evidence for complementarity in the dry compared to the wet seasons. While it appeared as though the simplicial complexes had more holes of various dimensions during the dry seasons, we were unable to detect statistically significant differences using a composite index of filtration complementarity. During the dry seasons, spider monkeys tend to rely on less abundant tree species and thus exploit a more heterogenous environment. During the wet season, the hyperabundance of Brosimum alicastrum [20] would make it less valuable to share information about scarce sources of food. However, we did not find significant differences in the temporal variation of the biweekly fruit abundance between the seasons. While our sampled number of seasons is small, it is possible that rather than temporal variation, it is the spatial variation that is higher in the dry than in the wet seasons. During the dry seasons, fruit is more scarce and dispersed, subgroups tend to be smaller and home ranges larger [20]. It would be interesting to test the information content of different regions of the home range, using more accurate tree distribution and abundance data than was available for this study.
@@
More generally, finding resources in a heterogenous, dynamic environment can be seen as a problem of pattern recognition, which can be solved effectively by systems with some degree of distributed information processing. This capacity can be found in various natural systems, recently described as ‘liquid brains’ [38], ranging from social insect colonies to the immune system. These systems process information via interactions between mobile agents that interact only temporarily when they happen to be within reach. As an example, Vining et al. [39] modelled a set of simple agents that could only communicate with others within a short radius but were able to move about, thus increasing the number of agents with whom they could exchange information over time. This system was able to solve a global problem of density classification (converging to the binary state that was more common in the beginning of the run amongst all agents). Crucially, agents were able to solve the most difficult versions of the problem (where the difference between the majority and minority binary state was very small) only when moving at intermediate speeds. This allowed them to balance the frequency of interactions with others in their current neighbourhood and those in other locations. In other words, an optimal degree of mixing between agents led the group to solve more difficult problems effectively [39]. This optimal degree of mixing is analogous to the optimization of $w$ we used here: an optimal area of intersection between individual use areas allows monkeys to coincide to share information while exploring enough area on their own which can be shared. Unique information, gathered by individual members of the group or subsets of them, becomes complementary when monkeys coincide in redundant areas and can share what they uniquely know. The redundancy/uniqueness balance is also more generally related to that between exploitation, where more redundant information is gathered about the environment, and exploration, where more unique information is gathered [40, 41].
@@
Collective information processing can result from individual members sharing complementary information, as we have found here, or it may also include synergistic information [42], whereby two or more individual estimations of a target variable not only are complementary in their estimation but combine their information in such a way as to produce new knowledge [2]. An example would be if one subset of individuals would contribute the location of a food source and another subset the timing of the fruiting of that source. The latter could be done by individuals who happened to visit the same trees repeatedly, gathering unique information about their fruiting status (as has been observed in other high fission-fusion species like the chimpanzee (Pan troglodytes; [43]). The resulting, combined knowledge by both subsets of individuals would be synergistic in the sense of allowing all of them to exploit the food source according to its location and timing. Further empirical and modelling work could explore these ideas and search for synergy in collective information processing about the foraging environment in this and other species.
@@
Using a theoretical prediction of the optimal balance between the use of redundant and unique areas, as well as simplicial complexes to represent higher-order interactions, we have been able to show complementarity in information sharing, a hallmark of collective information processing [2, 3]. This work exemplifies the usefulness of higher-order methods, as we have been able to uncover details in the spatial structure which could not be found in a dyadic representation of the same structure. While these methods come with greater complexity over the dyadic approach, they are extremely well suited to tackling problems surrounding collective behaviour. Spider monkeys with different knowledge of their foraging area contribute with different pieces of information to group-level knowledge, in such a way that the group as a whole can access more foraging trees in a dynamic landscape than any individual could on its own. This complementary information sharing in higher-order spatial networks would be a compelling example of collective intelligence in natural conditions [44].
@@ Methods
Ranging and subgroup composition data was collected by experienced observers between January 2012 and December 2017 in a habituated group of Geoffroy’s spider monkeys (Ateles geoffroyi) living around the Punta Laguna lake in the Otoch Ma’ax yetel Kooh protected area, in the Yucatan peninsula, Mexico. This group was the subject of a long-term, continuous study from 1997 to 2020 [45]. We defined dry seasons from November to April and wet seasons from May to October [20]. The spider monkeys’ foraging environment differs crucially between the two seasons, due mainly to a very abundant species (Brosimum alicastrum) fruiting during the wet season [20]. Therefore, fruit during the wet season is more abundant and homogeneously distributed, while during the dry season spider monkeys use more species that are less abundant overall and thus face a more heterogenous and uncertain environment [20].
We tracked changes in fruit abundance by recording, every two weeks, the number of trees with fruit for each species in a phenological trail. This trail contained a sample of 10 individual trees from 16 species, including 12 which comprise > 50% of the monkey’s diet [46, 47]. We estimated the density of each of these species using data collected in eight 0.25ha plots (total of 2ha) and fourty-eight 100m x 2m line transects (total 0.96 ha) within the home range of the spider monkeys (see [48] for more details about the vegetation sampling).
@@
We used individual locations for the estimation of individual core ranges, defined as the 60% utilization distribution generated with the adaptive Local Convex Hull (a-LoCoH) method developed by [49], using the T-LoCoH package [50] for the R programming environment [49] with a constant value of 15000 for the a parameter for hull construction across seasons. The use of the 60% utilization distribution as a meaningful area is validated by previous work where we showed that it lied at the inflection point in the curve of isopleth values versus proportion of home range covered [51]. Core ranges defined in this manner represent the area where individuals concentrated their regular movements, excluding more sporadically used areas. For more details on the calculation of individual core ranges, see [20].
@@
We calculated the intersections and unions between individual core ranges for each season using the Spatstat package [52] for the R programming environment [53]. For a set $C=\{C_{1},\dots,C_{n}\}$ of $n$ individual core ranges, we used the ratio denoted as $w\in[0,1]$ of the areas of their intersection and their union:
@@
We used the fruit abundance data from the phenological trail, together with the diameter at breast height of the trees monitored and the density for each species in the study site, to estimate fruit abundance and its variability in each season. We used the following formulation, based on [20]:
@@
where $\operatorname{IFA}_{f}\geq 0$ is the index of fruit abundance for the monitoring period $f$ based on the 16 species monitored, $\operatorname{Tf}_{i}$ is the proportion of trees with fruit of species $i$, $\operatorname{DBH}_{i}$ is the sum of the diameter at breast height over all the sampled trees with fruit of species $i$ during period $f$ and $\operatorname{Den}_{i}$ is the tree density for species $i$ in the study site. We considered the $\operatorname{IFA}_{f}$ values for all the monitoring periods in a season to obtain the mean and the coefficient of variation of the $\operatorname{IFA}_{f}$ per season.
@@
For each dataset (i.e., each season), we represented the entire space-sharing network of the study group as a simplicial complex $K$. The nodes of the network were all the adult individuals in the group during the particular season of the dataset (Table SI1). Since all core ranges coincided in some small common area (e.g. for sleep), we defined a maximal simplex size, $n_{\text{max}}$. This maximal size was needed because including interactions among more than $n_{\text{max}}$ individuals yielded a trivial simplicial complex structure (containing every possible simplex), and these larger interactions were less likely to be related to foraging activity, but rather to common sleeping areas in the centre of the group’s home range. Then there exists a simplex connecting a subset of individuals if their corresponding intersection/union ratio, $w$, is non-zero, and if the size of the subset is no more than $n_{\text{max}}$. Upon observation, the value of $n_{\text{max}}=6$ was the largest which consistently resulted in non-trivial structures across all seasons. We computed the Betti numbers of the simplicial complexes for all seasons.
@@
We investigated how the final simplicial complex, $K$, was formed when adding simplices in decreasing order of redundancy. Between subsets of the same size, we assumed a higher relative spatial overlap, quantified by $w$, corresponded to a more redundant interaction. However, what constituted a redundant spatial overlap for information transfer depends upon the number of individuals in the subset, and therefore we required a scaling of the $w$ values with their corresponding subset size in order to make fair comparisons. To this end, we computed the $w$ value from the spatial overlap of $n$ individuals which optimised the information transfer when these individuals had equal domain knowledge and utilised their core ranges uniformly, balancing the size of the unique areas they could share information about with the opportunity to share them when coinciding with others. When this balance of redundant and unique information leads to an optimal information transfer, we define it as complementary. As detailed in the Supplementary Information, for $n$ individuals this optimal overlap was computed as:
@@
This expression describes the amount of overlap required for a group of $n$ individuals to maximise their transfer of information under the given assumptions of uniformity and independence of movements. We then defined a parameter $\alpha$ to describe how far from $w^{*}$ a given overlap was. For the $i$-th non-zero spatial overlap in the data, consisting of $n(i)$ many individuals and with relative overlap value $w_{i}$, the corresponding $\alpha$ value was given as the solution to:
@@
where $f(\alpha)$ is some arbitrary positive and decreasing function. We picked $f(\alpha)=5-\alpha$ as a linear function such that the corresponding $\alpha$ value for each set was non-zero, and each simplex has $\alpha>0$, thus ensuring we consistently begin our filtration with a trivial structure consisting only of 0-simplices. We varied $\alpha$ over the interval $[0,5)$. For $\alpha\leq 5$ each possible observation (including those with zero overlap, due to the inclusion of $\alpha=5$) is represented in the simplicial structure. Variation in the choice of the function $f$ would influence the axis of the persistence diagrams, but the same topological features would be observed.
Then, we defined a sequence of simplicial complexes $K_{0}\subseteq K_{1}\subseteq\dots\subseteq K_{N}=K$. Here, $K_{0}$ is the set of individuals in the study group (without any simplices connecting them) and each $K_{j}$ is a simplicial complex with those simplices with $\alpha$ below the corresponding $j$-th lowest $\alpha$ value, i.e. corresponding to the intersection/union ratio $w$ above the corresponding $w$ threshold (possibly adding additional lower-order simplices to satisfy downward closure). This last procedure is justified on the basis of the clear tendency of overlaps to occur toward the centre of the group’s range (see Figure 2), which makes it unlikely that the areas of $n$ individuals will overlap significantly but the overlaps between subsets of $n$ will not. More generally, in higher-order spatial studies, it is standard to assume that if a higher-order interaction is significant, then all lower-order sub-interactions should be considered significant [25]. This sequence is called a filtered simplicial complex, and $\alpha$ is called the filtration parameter [32]. Simplicial complexes which are lower in the sequence will consist of more redundant interactions in terms of their information sharing. We computed the Betti numbers for each simplicial complex in the sequence (the persistent homology) for all available years and seasons in order to examine how the higher-order structure varied with the redundancy of interactions. Analysis was performed using the Gudhi package for Topological Data Analysis, implemented in Python [54].
@@
We quantified how much evidence of complementarity in information sharing there was in a filtered simplicial complex through the filtration complementarity index. Structures with greater evidence of complementary sharing would be those with a larger number of higher order features (i.e. larger Betti numbers of high order), which persisted for longer throughout the filtration. Simplicial complexes with holes of higher dimensions may be described as more complex and present more evidence of complementary information sharing among a wider variety of simplices. We therefore defined the filtration complementarity index for season $i$ as:
@@
Here, $B_{i}^{(d)}(\alpha)$ is the $d$-th Betti number at the filtration value of $\alpha$ in the $i$-th season, $n_{i}$ is the number of individuals in the data in this season, and $\langle\cdot\rangle_{[0,5)}$ is the average operator over the filtration interval $[0,5)$. We scaled the number of connected components by the number of individuals as to not give seasons with larger sample sizes a higher score automatically (as the number of connected components is highly dependent upon $n_{i}$). We did not scale the higher-order features as we did not expect them to be (directly) impacted by sample size differences observed in our data. Higher-order features were given greater importance through the scaling by $d+1$. We also tested other definitions of the filtration complementary index, including non-weighted or which did not consider the connected components. All of them yielded similar results.
@@
We additionally studied which subsets were the most central in the simplicial complex across years and seasons. For this, we implemented the maximal simplicial degree centrality measure as defined in [31]. This is a generalisation of the standard degree centrality used in dyadic networks to simplicial complexes, which identifies those subsets that appear in the largest number of simplices. This metric is designed so that the downward closure assumption does not produce biases towards higher or lower order simplices, which are not automatically classed as more or less central just for being part of a nested structure. We computed the centrality value for each simplex in the simplicial complex for each year and season at a filtration value of $\alpha=4$ (meaning we do not consider subsets with more unique interactions). Other values of $\alpha$ yielded similar results. We analysed how simplicial centrality varied with simplex characteristics (size, average age, proportion of males and proportion of immigrants). When analysing how simplicial centrality varies with simplex size, we considered variation in the parameter $n_{\text{max}}\in\{4,5,6,7,8,9\}$ to ensure our results were qualitatively robust to variations in the distribution of simplex sizes.
@@
The individual core ranges analysed, as well as the data used for constructing the simplicial complexes, are publicly available as part of the repository cited above.